[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jul 16, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Jul 17 03:08:22 UTC 2016


HSU Shift Log for Sat Jul 16 2016 23:08:16 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: TAFB fixes were very late to import into ATCF (at least 30 min  
after submitted). Also here's a bug that I think has been recorded: When  
the intensity is modified at T+3, it results in the TC status reverting to  
LO/DB, and messes up your product headers.

N-AWIPS/AWIPS2 issues: none

Other Techincal issues: none
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160717/74bf0c01/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list